OpenStudy (anonymous):

how do i divide 7sqrt of 2 by the sqrt of 2?

OpenStudy (anonymous):

\[7\sqrt{2}\div \sqrt{2}\]

OpenStudy (anonymous):

What does \(7\sqrt{2}\) mean?

OpenStudy (anonymous):

it means \[7\sqrt{2}\]

OpenStudy (anonymous):

In english. I know what it means. I want to make sure you know what it means.

OpenStudy (anonymous):

i do, i just need to figure out how to divide it by the sqrt of 2

OpenStudy (anonymous):

If you knew what it meant I suspect you wouldn't have any trouble. What is being done to the 7 and the \(\sqrt{2}\) in the expression \(7\sqrt{2}\)

OpenStudy (anonymous):

never mind, the answer is 7 right?

OpenStudy (anonymous):

indeed! =)

OpenStudy (anonymous):

It's important to look beyond the symbols and remind yourself what it means =)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Very welcome =)